A magnetic field increases from 0 to 0.25T in 1.8s

Part A

How many turns of wire are needed in a circular coil 12cm in diameter to produce an induced emf of 6.3V ?
Answer #1

induced emf= NA dB/dt
Area A= pi*r^2
   =pi*0.06^2
=0.01131 m^2

dB/dt =0.25 /1.8=0.1389 T/sec

induced emf= NA dB/dt
6.3=N*0.01131*0.1389
N=4010

Number of turns=4010